# Data-driven Error Estimation: Upper Bounding Multiple Errors without Class Complexity as Input

### v3

@@ 4 Conclusion
$\Pr\left(e_{h^{*}}\leq\max_{h\in\mathcal{G}(S_{\text{err}})}\hat{u}(S_{\text{err}},h,\delta)|\mathcal{E}(s)\right)\stackrel{{\scriptstyle(i)}}{{\geq}}\Pr_{S{\text{err}}\sim\mathcal{D}}(\mathcal{W}^{*}(S_{\text{err}})|\mathcal{E}(s))\stackrel{{\scriptstyle(ii)}}{{=}}\Pr_{S{\text{err}}\sim\mathcal{D}}\Big{(}\hat{u}(S_{\text{err}},h^{*},\delta)\geq e_{h^{*}}\Big{|}\mathcal{E}(s)\Big{)}\stackrel{{\scriptstyle(iii)}}{{\geq}}1-\delta$ (18)
Here (i) holds by (17), (ii) holds by the definition of $\mathcal{W}^{*}(S_{\text{err}})$, and (iii) holds by 2.4. Furthermore, if we can construct a sequence of classes $\{\mathcal{H}_{k}(S_{\text{err}})\}_{k=0}^{\infty}\subseteq\mathcal{H}$ where for each $k$, $\mathcal{W}^{*}(S_{\text{err}})\subseteq\{h^{*}\in\mathcal{H}_{k}(S_{\text{err}})\}$, we can consider $\mathcal{H}_{\infty}$, and apply (18) 2.4, $\Pr\left(e_{h^{*}}\leq\max_{h\in\cap_{k}\mathcal{H}_{k}(S_{\text{err}})}\hat{u}(S_{\text{err}},h,\delta)\right)\geq 1-\delta$.
@@
We claim $\mathcal{W}^{*}(S_{\text{err}})\subseteq\{h^{*}\in\mathcal{H}_{k}(S_{\text{err}})\}$ for all $k\in\mathbb{N}$. To show this we assume $\mathcal{W}^{*}(S_{\text{err}})$ holds and deduce $h^{*}\in\mathcal{H}_{k}(S_{\text{err}})\forall k$ by induction. First note that this holds true for $k=0$ since $h^{*}\in\mathcal{H}_{0}(S_{\text{err}})=\mathcal{H}$. Now, assume $h^{*}\in\mathcal{H}_{k}(S_{\text{err}})$. Recall that $e_{h^{*}}=\theta_{h^{*}}-\hat{\theta}_{h^{*}}\geq c-\hat{\theta}_{h^{*}}$ by 2.3. Thus, by definition of $\mathcal{H}_{k+1}(S_{\text{err}})$, we have $h^{*}\in\mathcal{H}_{k+1}(S_{\text{err}})$. Thus we have arrived at the following localization corollary.
Appendix B Additional details for Section 3.2
In this section, we show how to formally convert the example in Section 1 into the setting described in Section 3.2. Recall the goal of constructing simultaneously valid CIs on average test scores for subgroups (e.g.“female students”, “fourth-graders”, etc.) indexed by $\mathcal{H}$ [^18] With mild abuse of notation, we let any $h\in\mathcal{H}$ denote both the subgroup index and subgroup itself. More formally, let $(X,T)\in\mathcal{X}\times[-1,1]$ be a pair of random variables corresponding to a randomly sampled student’s feature vector and their test score respectively. The average test score for any subgroup $h$ is then given by $\theta_{h}$ in (20).
$\theta_{h}:=\mathop{\mathbb{E}}[T|X\in h]=\mathop{\mathbb{E}}[Y(h)],$ (20)
$Y(h):=\frac{T\cdot\mathbf{1}(X\in h)}{p_{h}},\;p_{h}:=\Pr(X\in h).$
For ease of exposition, we assume the probability ($p_{h}$) of a random student belonging to subgroup $h$ is known to us and is lower bounded by $1/M$ for some $M\in[1,\infty)$. [^19] We can always estimate $p_{h}$ from $S_{\text{def}}$ and reset $\mathcal{H}\leftarrow\{h\in\mathcal{H}|p_{h}\geq 1/M\}$Corollary 3.3 would still hold. The remarks would also hold if cross-fitting (Newey & Robins, 2018) were used to account for unknown $p_{h}$ in estimates from $S_{\text{def}}$ and $S_{\text{err}}$.
@@
Remark B.1 (Asymptotic Distribution of $\hat{u}$).
@@
B.1 Details for Figures
  • Figure 1 is generated as follows:
    We generate $n=10000$ i.i.d. samples from a $\mathcal{N}(0,1)$ distribution, and empirically estimate the quantiles of this distribution. Next, we define our errors to be the absolute value of the difference between these estimates and the true theoretical quantiles, normalized so that each error is $|\mathcal{N}(0,1)|$ (to do this, we rely on the distribution of a sample quantile, $\hat{\theta}_{q}\sim\mathcal{N}\left(\theta_{q},\frac{q(1-q)}{nf(\theta_{q})^{2}}\right)$. As we don’t assume access to the density function $f$, we estimate this using a gaussian kernel estimator. Once we have defined our errors, we use the error estimation theory above, along with an additional dataset of $n=10000$ points to create bounds on the maximum error. We compare this to the bounds generated by union bounding for the original dataset. We repeat each experiment 100 times and plot the average.
  • Figure 2 is generated as follows:
    We define a model class of Neural Networks, then use the Iris dataset from UCI ML Repo as our dataset. We first train a model $g^{*}$ on the dataset to serve as the true realizable model, so we alter the data to be generated by the model. We then add noise to the dataset, split it into $S_{\text{def}}$ and $S_{\text{err}}$, and train a secondary model $g_{est}$ on $S_{\text{def}}$. Next we evaluate the risk of model $g_{\text{def}}$ over increasingly noisy versions of $S_{\text{def}}$ and $S_{\text{err}}$, and use the error estimation theory to create bounds on excess risk. To evaluate the excess risk bound, we iteratively search for a model that maximizes the difference in empirical risk under $S_{\text{def}}$ and $S_{\text{err}}$, subject to an increasingly strict localization constraint. Additionally, instead of relying on Hoeffding’s inequality we use conformal methods to estimate a high probability upper bound on a single error.
Appendix C Multiple Hypothesis Testing
@@
[…]
Corollary C.1.
@@
Note that FWER for Corollary C.1 is at most $\delta$ for any choice of $w$. Now let us briefly discuss the choices for $w=(\iota,b)$. A simple choice is to let $b\equiv\iota\equiv 1$, this choice makes Corollary C.1 a direct application of Corollary 3.3. However, while keeping FWER less than $\delta$, a more intelligent choice for $w$ will allow us to discover more $h$ for which the alternative hypothesis is true (larger $\mathcal{H}_{r}$). Note that even for rejecting the single null hypothesis $\theta_{h}=0$ with dataset $S_{\text{def}}$, we would require $\hat{\theta}_{h}>\hat{\sigma}_{\text{def},h}z_{\delta}/\sqrt{n}$. Hence for multiple testing, we only need to focus our attention on $h$ satisfying this condition. It would thus be reasonable to choose $\iota(h)=\mathbf{1}(\hat{\theta}_{h}>\hat{\sigma}_{\text{def},h}z_{\delta}/\sqrt{n})$, ensuring $\hat{\xi}_{w}(S_{\text{err}},\delta)$ is not larger than necessary. Using $b(h)=\hat{\sigma}_{\text{def},h}\min(1,1/|\hat{\theta}_{h}|)$ is also a reasonable choice – this makes the tests in (22) more comparable – allowing for $\mathcal{H}_{r}$ to be much larger when possible. If we are willing to incur a larger FWER in exchange for more nulls rejected, we can set $\iota(h)$ as zero more often and extrapolate the test in (22) to such $h$. Such approaches could allow us to tackle objectives like controlling the false discovery rate (FDR). Analyzing the choices of $w$ and its various implications is beyond the scope of this paper.
Appendix D Contextual Bandit Pipeline with Error Estimation
@@
$R_{f}(p):=\mathop{\mathbb{E}}_{(x,r)\sim D}\mathop{\mathbb{E}}_{a\sim p(\cdot|x)}[f(x,a)],\quad\text{and }V(p,\pi):=\mathop{\mathbb{E}}_{(x,r)\sim D,a\sim\pi(\cdot|x)}\bigg{[}\frac{\pi(a|x)}{p(a|x)}\bigg{]}.$ (25)
@@
Objective. Contextual bandit algorithms run in epochs (batches). At the start of each epoch, a new exploration (data-collection) policy is computed and is used to select arms for each round in that epoch. Consider an epoch that starts at round $\tau+1$. To construct an exploration policy $p$ for this epoch, we leverage dataset $S$ corresponding to the first $\tau$ rounds. Arguably, designing a contextual bandit algorithm boils down to designing an algorithm that uses data from prior epochs to construct the new exploration policy $p$. Hence, to simplify our analysis and focus on algorithmic insights, we study an objective that is solely a function of $p$. In particular, our objective is to construct $p$ in order to minimize the “optimal cover”, defined as $V(p,\pi^{*})$. The optimal cover is proportional to the variance of evaluating the optimal policy and has been shown to be an effective surrogate objective for pure exploration (see discussions in Krishnamurthy et al., 2023; Maurer & Pontil, 2009). [^24] At a high-level, minimizing the variance of evaluating the optimal policy allows variance-penalized policy learning algorithms to ensure better guarantees (Xu & Zeevi, 2020). Hence minimizing the optimal cover is seen as a surrogate objective for pure exploration. Furthermore, although there can be some trade-off between minimizing $V(p,\pi^{*})$ and maximizing $R(p)$, minimizing the optimal cover encourages behavior similar to the optimal policy and thus naturally results in exploration policies with high expected reward.
@@
D.1 Contextual Bandit Pipeline
@@
6:  Estimate $\hat{U}_{\text{con}},\leftarrow\text{CBConErrorEstimator}(S_{\text{def}},S_{\text{err}},\hat{f},\hat{\pi}_{\text{con}},\hat{R}_{\text{def}},\hat{\pi}_{\text{elim}},\delta,\alpha_{\text{err}})$ such that (29) holds.
@@
8:  Select risk adjustment parameter $\eta\in[1,K]$. {Attempting to minimizes the optimal cover upper bound in Theorem D.1.}
@@
[…]
@@
This section describes our contextual bandit pipeline as detailed in Algorithm 1. To facilitate a truly modular contextual bandit pipeline, we employ “black box” subroutines without specific requirements on output quality whenever possible. This allows for the use of a wide range of heuristic subroutines that perform well in practice but may lack strong theoretical guarantees (e.g., methods based on deep learning or fine-tuning). To create bounds on the suboptimality of the performance of these black box subroutines, we use the error estimation theory derived above. Therefore, our algorithm begins by splitting the input dataset $S$ into $S_{\text{def}}$ and $S_{\text{err}}$,262626ensuring that the rounds in $S_{\text{err}}$ follow those in $S_{\text{def}}$. and then processes this data using several subroutines. As a result of the doubling epoch schedule, the $S_{\text{err}}$ dataset is collected under a single exploration policy $p_{\text{err}}$. We take as input a bound on the optimal cover of this exploration policy, $\alpha_{\text{err}}\geq V(p_{\text{err}},\pi^{*})$. The algorithm will be using error estimation theory in two separate instances. Additionally, we set off $S_{\text{err},B}$ to bound the optimal cover of the next exploration policy. Accordingly we split $S_{\text{err}}$ into $S_{\text{err},\text{elim}}$, $S_{\text{err},con}$, and $S_{\text{err},B}$. (see Figure 3 for outline of data allocation in round $\tau$).
@@
To estimate a high-probability upper bound on the $\hat{R}_{\text{elim}}(\hat{\pi}_{\text{elim}})-\hat{R}_{\text{elim}}(\pi^{*})$ from the output of BlackBoxPolicyEvaluator, we use a subroutine (CBElimErrorEstimator) on the dataset $S_{\text{err},\text{elim}}$. We require the output from this subroutine ($\hat{U}_{\text{elim}}$) to satisfy (27) and discuss one approach (based on error estimation) to designing such a subroutine in Section D.2 using $S_{\text{err},\text{elim}}$. With this error bound, we use subroutine (ArmEliminator) to construct $\hat{g}:\mathcal{X}\rightarrow 2^{\mathcal{A}}$ satisfying (28). Note that (28) is not a statistical requirement, and the challenge in designing such a subroutine is purely computational. A trivial way to satisfy (28) is to choose $\hat{g}(x)=\mathcal{A}$ for all $x$. We describe a more non-trivial and implementable approach to designing such a subroutine in Section D.3.
@@
Once again, we describe an approach using error estimation to estimate a high probability upper bound on $R_{\hat{f}}(\hat{\pi}_{\text{con}})-R_{\hat{f}}(\pi^{*})$, a measure of the quality of $\hat{f}$ as the output of BlackBoxRegressor. We require the output ($\hat{U}_{\text{con}}$) to satisfy (29) and design a subroutine (CBConErrorEstimator) in Section D.2 using data $S_{\text{err},\text{con}}$.
@@
Finally, we can upper bound both the optimal cover and the maximum inverse probability of sampling the arm recommended by the optimal policy (see (31)). Here we use $S_{\text{err},B}$. See Theorem D.1 for formal theoretical guarantees.
@@
$V(p,\pi^{*})\leq\alpha_{\tau+1}=\mathop{\mathbb{E}}_{x\sim S_{\text{err},B}}\Bigg{[}\frac{|C(x,\beta_{\max}/\eta)|}{1-\beta_{\max}+\beta_{\max}\frac{|C(x,\beta_{\max}/\eta)|}{|\hat{g}(x)|}}\Bigg{]}+\frac{\max_{x}|\hat{g}(x)|}{\eta}+\sqrt{\frac{\max_{x}|\hat{g}(x)|\log(3/\delta)}{\beta_{\max}|S_{\text{err},B}|}}.$
Note that with an appropriate choice of $\eta$, the optimal cover bound from Theorem D.1 primarily depends on the average size of the set $C(x,\beta_{\max}/\eta)$ – which would be small when the CBErrorEstimator outputs ($\hat{U}_{\text{elim}}$ and $\hat{U}_{\text{con}}$) are small – which in turn can only happen when outputs of the black box subroutines have fundamentally small errors.
D.2 Error Estimation for Contextual Bandits
In this section, we apply error estimation theory to develop the two CBErrorEstimator subroutines, ensuring they meet the criteria specified in equations (27) and (29). We present an overview of the terms used in Table 4. While CBElimErrorEstimatorimplements error estimation in a manner very similar to the application in Section 3.3, CBConErrorEstimatorinvolves an extra step, as the ”estimate” quantity is not directly accessible. We outline the methods and algorithms in this section, and provide detailed proofs in Section D.4.2.

### v4

@@ D.1 Replacing $\max$ with $\sup$ for infinite $\mathcal{H}$
$\Pr\left(e_{h^{*}}\leq\max_{h\in\mathcal{G}(S_{\text{err}})}\hat{u}(S_{\text{err}},h,\delta)|\mathcal{E}(s)\right)$ $\stackrel{{\scriptstyle(i)}}{{\geq}}\Pr_{S{\text{err}}\sim\mathcal{D}}(\mathcal{W}^{*}(S_{\text{err}})|\mathcal{E}(s))$
$\stackrel{{\scriptstyle(ii)}}{{=}}\Pr_{S{\text{err}}\sim\mathcal{D}}\Big(\hat{u}(S_{\text{err}},h^{*},\delta)\geq e_{h^{*}}\Big|\mathcal{E}(s)\Big)\stackrel{{\scriptstyle(iii)}}{{\geq}}1-\delta$
Here (i) holds by (21), (ii) holds by the definition of $\mathcal{W}^{*}(S_{\text{err}})$, and (iii) holds by 2.4. Furthermore, if we can construct a sequence of classes $\{\mathcal{H}_{k}(S_{\text{err}})\}_{k=0}^{\infty}\subseteq\mathcal{H}$ where for each $k$, $\mathcal{W}^{*}(S_{\text{err}})\subseteq\{h^{*}\in\mathcal{H}_{k}(S_{\text{err}})\}$, we can consider $\mathcal{H}_{\infty}$, and by 2.4, $\Pr\left(e_{h^{*}}\leq\max_{h\in\cap_{k}\mathcal{H}_{k}(S_{\text{err}})}\hat{u}(S_{\text{err}},h,\delta)\right)\geq 1-\delta$.
@@
We claim $\mathcal{W}^{*}(S_{\text{err}})\subseteq\{h^{*}\in\mathcal{H}_{k}(S_{\text{err}})\}$ for all $k\in\mathbb{N}$. To show this we assume $\mathcal{W}^{*}(S_{\text{err}})$ holds and deduce $h^{*}\in\mathcal{H}_{k}(S_{\text{err}})\forall k$ by induction. First note that this holds true for $k=0$ since $h^{*}\in\mathcal{H}_{0}(S_{\text{err}})=\mathcal{H}$. Now, assume $h^{*}\in\mathcal{H}_{k}(S_{\text{err}})$. Recall that $e_{h^{*}}=\theta_{h^{*}}-\hat{\theta}_{h^{*}}\geq c-\hat{\theta}_{h^{*}}$ by 2.3. Thus, by definition of $\mathcal{H}_{k+1}(S_{\text{err}})$, we have $h^{*}\in\mathcal{H}_{k+1}(S_{\text{err}})$.
Appendix E Additional details for Section 3.2
In this section, we show how to formally convert the example in Section 1 into the setting described in Section 3.2. Recall the goal of constructing simultaneously valid CIs on average test scores for subgroups (e.g.“female students”, “fourth-graders”, etc.) indexed by $\mathcal{H}$. [^21] With mild abuse of notation, we let any $h\in\mathcal{H}$ denote both the subgroup index and subgroup itself. More formally, let $(X,T)\in\mathcal{X}\times[-1,1]$ be a pair of random variables corresponding to a randomly sampled student’s feature vector and their test score respectively. The average test score for any subgroup $h$ is then given by $\theta_{h}$ in (23).
$\theta_{h}=\mathop{\mathbb{E}}[T|X\in h]=\mathop{\mathbb{E}}[Y(h)],$ (23)
$Y(h)=\frac{T\cdot\mathbf{1}(X\in h)}{p_{h}},\;p_{h}=\Pr(X\in h).$
For ease of exposition, we assume the probability ($p_{h}$) of a random student belonging to subgroup $h$ is known to us and is lower bounded by $1/M$ for some $M\in[1,\infty)$. [^22] We can always estimate $p_{h}$ from $S_{\text{def}}$ and reset $\mathcal{H}\leftarrow\{h\in\mathcal{H}|p_{h}\geq 1/M\}$Corollary 3.4 would still hold. The remarks would also hold if cross-fitting (Newey and Robins, 2018) were used to account for unknown $p_{h}$ in estimates from $S_{\text{def}}$ and $S_{\text{err}}$.
@@
Remark E.1 (Asymptotic Distribution of $\hat{u}$).
@@
Appendix F Multiple Hypothesis Testing
@@
We convert the setup from Section 3.2 into a multiple-hypothesis testing problem. Suppose, for every $h\in\mathcal{H}$, the null hypothesis is $\theta_{h}=0$ [^25] Without loss of generality, the threshold can be zero (by appropriately adjusting the estimand). and the alternative hypothesis is $\theta_{h}>0$. We reuse $\hat{u}$ constructed in (10), and introduce additional weight functions $\iota$ and $b$ over the class of hypotheses. Here $\iota$ is an indicator function allowing us to select which hypotheses to include in the testing procedure, and $b$ is a weighting function that allows us to adjust the relative importance for rejecting certain hypotheses in the error estimation process. These functions introduce flexibility in our approach, allowing us to use prior knowledge to focus on a specific subset or emphasize certain hypotheses.
Corollary F.1.
@@
Note that FWER for Corollary F.1 is at most $\delta$ for any choice of $w$. Now let us briefly discuss the choices for $w=(\iota,b)$. A simple choice is to let $b\equiv\iota\equiv 1$, this choice makes Corollary F.1 a direct application of Corollary 3.4. However, while keeping FWER less than $\delta$, a more intelligent choice for $w$ will allow us to discover more $h$ for which the alternative hypothesis is true (larger $\mathcal{H}_{r}$). Note that even for rejecting the single null hypothesis $\theta_{h}=0$ with dataset $S_{\text{def}}$, we would require $\hat{\theta}_{h}>\hat{\sigma}_{\text{def},h}z_{\delta}/\sqrt{n}$. Hence for multiple testing, we only need to focus our attention on $h$ satisfying this condition. It would thus be reasonable to choose $\iota(h)=\mathbf{1}(\hat{\theta}_{h}>\hat{\sigma}_{\text{def},h}z_{\delta}/\sqrt{n})$, ensuring $\hat{\xi}_{w}(S_{\text{err}},\delta)$ is not larger than necessary. Using $b(h)=\hat{\sigma}_{\text{def},h}\min(1,1/|\hat{\theta}_{h}|)$ is also a reasonable choice – this makes the tests in (25) more comparable – allowing for $\mathcal{H}_{r}$ to be much larger when possible. If we are willing to incur a larger FWER in exchange for more nulls rejected, we can set $\iota(h)$ as zero more often and extrapolate the test in (25) to such $h$. Such approaches could allow us to tackle objectives like controlling the false discovery rate (FDR). Analyzing the choices of $w$ and its various implications is beyond the scope of this paper.
Appendix G Improved performance via cross-fitting
Our basic error-estimation construction uses a single split: a defining dataset $S_{\text{def}}$ to form the estimates $\{\hat{\theta}_{h}\}_{h\in\mathcal{H}}$ (and the induced class $\mathcal{E}(s)$), and an independent error-estimation dataset $S_{\text{err}}$ to compute the surrogate bounds $\hat{u}(S_{\text{err}},h,\delta)$. This split can inflate the resulting uniform bound, since each direction uses only half the data for estimation.
A simple mitigation is two-way dataset switching (a K=2K{=}2 cross-fitting variant). Let the full sample be split into two folds $S^{(1)}$ and $S^{(2)}$ (e.g., equal size). We run the method in both directions: (i) define $\{\hat{\theta}_{h}\}$ using $S^{(1)}$ and compute the uniform bound using $S^{(2)}$, yielding
$\hat{\xi}^{1\to 2}(\lambda)\;:=\;\sup_{h\in\mathcal{H}}\hat{u}\!\left(S^{(2)},h,\lambda\right);$
(ii) swap the roles of the folds, yielding
$\hat{\xi}^{2\to 1}(\lambda)\;:=\;\sup_{h\in\mathcal{H}}\hat{u}\!\left(S^{(1)},h,\lambda\right).$
We then report the switched bound
$\hat{\xi}^{\min}(\delta)\;:=\;\min\Big\{\hat{\xi}^{1\to 2}(\delta/2),\;\hat{\xi}^{2\to 1}(\delta/2)\Big\}.$
Figure 5: Effect of two-way dataset switching in the correlated finite-class simulation (100 runs per $\alpha$). Orange: error-estimation bound using the original split direction; dashed orange: the same procedure after swapping the roles of $S_{\text{def}}$ and $S_{\text{err}}$; red: the switched bound $\hat{\xi}^{\min}(\delta)$. Green curves show the corresponding realized max errors under each direction, and the blue curve shows the max error computed without a split (oracle baseline). Switching and taking the minimum substantially tightens the bound, often mitigating a large fraction of the split-induced looseness.
Empirically, Figure 5 shows three regimes: (i) the original direction, (ii) the switched direction, and (iii) taking the minimum of the two valid bounds (with $\delta/2$ in each direction). In this simulation, the min-of-two procedure often reduces conservatism by a sizable constant factor, bringing the bound closer to the no-split oracle baseline.
Beyond two folds.
More generally, one can repeat this idea over multiple random resplits (or $K$ folds), compute $m$ valid bounds at confidence levels $\delta/m$, and take their intersection (equivalently, the minimum of the $m$ upper bounds). At some point the additional Bonferroni cost from using $\delta/m$ will outweigh the benefit of intersecting, so selecting the number of resplits is a practical tradeoff; understanding this tradeoff more sharply is an interesting direction for future work.
Appendix H Contextual Bandit Pipeline with Error Estimation
@@
$R_{f}(p)=\mathop{\mathbb{E}}_{(x,r)\sim D}\mathop{\mathbb{E}}_{a\sim p(\cdot|x)}[f(x,a)],\quad\text{and }V(p,\pi)=\mathop{\mathbb{E}}_{(x,r)\sim D,a\sim\pi(\cdot|x)}\bigg[\frac{\pi(a|x)}{p(a|x)}\bigg].$ (28)
@@
Objective. Contextual bandit algorithms run in epochs (batches). At the start of each epoch, a new exploration (data-collection) policy is computed and is used to select arms for each round in that epoch. Consider an epoch that starts at round $\tau+1$. To construct an exploration policy $p$ for this epoch, we leverage dataset $S$ corresponding to the first $\tau$ rounds. Arguably, designing a contextual bandit algorithm boils down to designing an algorithm that uses data from prior epochs to construct the new exploration policy $p$. Hence, to simplify our analysis and focus on algorithmic insights, we study an objective that is solely a function of $p$. In particular, our objective is to construct $p$ in order to minimize the “optimal cover”, defined as $V(p,\pi^{*})$. The optimal cover is proportional to the variance of evaluating the optimal policy and has been shown to be an effective surrogate objective for pure exploration (see discussions in Krishnamurthy et al., 2023; Maurer and Pontil, 2009). [^27] At a high-level, minimizing the variance of evaluating the optimal policy allows variance-penalized policy learning algorithms to ensure better guarantees (Xu and Zeevi, 2020). Hence minimizing the optimal cover is seen as a surrogate objective for pure exploration. Furthermore, although there can be some trade-off between minimizing $V(p,\pi^{*})$ and maximizing $R(p)$, minimizing the optimal cover encourages behavior similar to the optimal policy and thus naturally results in exploration policies with high expected reward.
@@
H.1 Contextual Bandit Pipeline
@@
6: Estimate $\hat{U}_{\text{con}},\leftarrow\text{CBConErrorEstimator}((S_{\text{def}},S_{\text{err},\text{elim}}),S_{\text{err},\text{con}},\hat{f},\hat{\pi}_{\text{con}},\hat{R}_{\text{def}},\hat{\pi}_{\text{elim}},\delta,\alpha_{\text{err}})$ such that (32) holds.
@@
8: Select risk adjustment parameter $\eta\in[1,K]$. {Attempting to minimizes the optimal cover upper bound in Theorem H.1.}
@@
$S_{\text{def}}$$S_{\text{err}}$Next round explored with $p_{\text{err}}$$S_{\text{err},\text{elim}}$$S_{\text{err},\text{con}}$$S_{\text{err},B}$
@@
This section describes our contextual bandit pipeline as detailed in Algorithm 1. To facilitate a truly modular contextual bandit pipeline, we employ “black box” subroutines without specific requirements on output quality whenever possible. This allows for the use of a wide range of heuristic subroutines that perform well in practice but may lack strong theoretical guarantees (e.g., methods based on deep learning or fine-tuning). To create bounds on the suboptimality of the performance of these black box subroutines, we use the error estimation theory derived above. Therefore, our algorithm begins by splitting the input dataset $S$ into $S_{\text{def}}$ and $S_{\text{err}}$,292929ensuring that the rounds in $S_{\text{err}}$ follow those in $S_{\text{def}}$. and then processes this data using several subroutines. As a result of the doubling epoch schedule, the $S_{\text{err}}$ dataset is collected under a single exploration policy $p_{\text{err}}$. We take as input a bound on the optimal cover of this exploration policy, $\alpha_{\text{err}}\geq V(p_{\text{err}},\pi^{*})$. The algorithm will be using error estimation theory in two separate instances. Additionally, we set off $S_{\text{err},B}$ to bound the optimal cover of the next exploration policy. Accordingly we split $S_{\text{err}}$ into $S_{\text{err},\text{elim}}$, $S_{\text{err},con}$, and $S_{\text{err},B}$. (see Figure 6 for outline of data allocation in round $\tau$).
@@
To estimate a high-probability upper bound on the $\hat{R}_{\text{elim}}(\hat{\pi}_{\text{elim}})-\hat{R}_{\text{elim}}(\pi^{*})$ from the output of BlackBoxPolicyEvaluator, we use a subroutine (CBElimErrorEstimator) on the dataset $S_{\text{err},\text{elim}}$. We require the output from this subroutine ($\hat{U}_{\text{elim}}$) to satisfy (30) and discuss one approach (based on error estimation) to designing such a subroutine in Section H.2 using $S_{\text{err},\text{elim}}$. With this error bound, we use subroutine (ArmEliminator) to construct $\hat{g}:\mathcal{X}\rightarrow 2^{\mathcal{A}}$ satisfying (31). Note that (31) is not a statistical requirement, and the challenge in designing such a subroutine is purely computational. A trivial way to satisfy (31) is to choose $\hat{g}(x)=\mathcal{A}$ for all $x$. We describe a more non-trivial and implementable approach to designing such a subroutine in Section H.3.
@@
Once again, we describe an approach using error estimation to estimate a high probability upper bound on $R_{\hat{f}}(\hat{\pi}_{\text{con}})-R_{\hat{f}}(\pi^{*})$, a measure of the quality of $\hat{f}$ as the output of BlackBoxRegressor. We require the output ($\hat{U}_{\text{con}}$) to satisfy (32) and design a subroutine (CBConErrorEstimator) in Section H.2 using data $S_{\text{err},\text{con}}$.
@@
Finally, we can upper bound both the optimal cover and the maximum inverse probability of sampling the arm recommended by the optimal policy (see (34)). Here we use $S_{\text{err},B}$. See Theorem H.1 for formal theoretical guarantees.
@@
$V(p_{\tau+1},\pi^{*})\leq\alpha_{\tau+1}=\mathop{\mathbb{E}}_{x\sim S_{\text{err},B}}\Bigg[\frac{|C(x,\beta_{\max}/\eta)|}{1-\beta_{\max}+\beta_{\max}\frac{|C(x,\beta_{\max}/\eta)|}{|\hat{g}(x)|}}\Bigg]+\frac{\max_{x}|\hat{g}(x)|}{\eta}+\sqrt{\frac{\max_{x}|\hat{g}(x)|\log(3/\delta)}{\beta_{\max}|S_{\text{err},B}|}}.$
Note that with an appropriate choice of $\eta$, the optimal cover bound from Theorem H.1 primarily depends on the average size of the set $C(x,\beta_{\max}/\eta)$ – which would be small when the CBErrorEstimator outputs ($\hat{U}_{\text{elim}}$ and $\hat{U}_{\text{con}}$) are small – which in turn can only happen when outputs of the black box subroutines have fundamentally small errors.
H.2 Error Estimation for Contextual Bandits
In this section, we apply error estimation theory to develop the two CBErrorEstimator subroutines, ensuring they meet the criteria specified in equations (30) and (32). We present an overview of the terms used in Table 4. While CBElimErrorEstimator implements error estimation in a manner very similar to the application in Section 3.3, CBConErrorEstimator involves an extra step, as the ”estimate” quantity is not directly accessible. We outline the methods and algorithms in this section, and provide detailed proofs in Section H.4.2.